Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufactu2. How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calcula3. What was the total manufacturing cost assigned to Job P? (Do not round intermediate calculations.) Total manufacturing cos4. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ5. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.) Total manufacturing cos6. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ7. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establi8. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold9. What were the companys predetermined overhead rates in the Molding Department and the Fabrication Department? (Round your10. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do n
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Sweeten
Answer 1
Plant wide predetermined OH Rate Molding Fabrication Total Note
Machine Hours        2,500.00        1,500.00     4,000.00 A
Variable OH Rate                3.00                3.80 B
Variable OH cost        7,500.00        5,700.00 13,200.00 C=A*B
Fixed Manufacturing overhead      14,000.00      17,400.00 31,400.00 D
Total OH 44,600.00 E=C+D
Plant wide predetermined OH Rate           11.15 F=E/A
Answer 2 Job P Job Q Note
Actual Machine hours        5,500.00        4,900.00 G
Plant wide predetermined OH Rate             11.15              11.15 See F
Overhead applied     61,325.00      54,635.00 H=G*F
Answer 3 Amount $ Note
Direct Materials      29,000.00
Direct Labor      33,800.00
Overhead applied      61,325.00 See H
Manufacturing cost assigned to Job P 124,125.00 I
Answer 4 Amount $
Manufacturing cost assigned to Job P 124,125.00 See I
Number of units             20.00 J
Unit product cost        6,206.25 K=I/J
Answer 5 Amount $
Direct Materials      16,000.00
Direct Labor      13,900.00
Overhead applied      54,635.00 See H
Manufacturing cost assigned to Job Q     84,535.00 L
Answer 6 Amount $
Manufacturing cost assigned to Job Q      84,535.00 See L
Number of units             30.00 M
Unit product cost        2,817.83 N=L/M
Answer 7 Job P Job Q Note
Manufacturing cost assigned 124,125.00      84,535.00 O=K,N
Mark up at 80%      99,300.00      67,628.00 P=O*80%
Total price for the job 223,425.00    152,163.00 Q=P+O
Number of units             20.00              30.00 R=J,M
Selling price per unit     11,171.25        5,072.10 S=Q/R
Answer 8 Amount $ Note
Manufacturing cost assigned to Job P 124,125.00 See K
Manufacturing cost assigned to Job Q      84,535.00 See N
Cost of goods sold 208,660.00
Answer 9
Calculation of predetermined OH Rate Molding Fabrication Note
Machine Hours        2,500.00        1,500.00 See A
Fixed Manufacturing overhead      14,000.00      17,400.00 See D
Fixed Manufacturing overhead rate                5.60              11.60 T=D/A
Variable OH Rate                3.00                3.80 See B
Departmental predetermined OH Rate                8.60              15.40 U=T+B
Answer 10- Job P Molding Fabrication Total Note
Machine hours used        3,300.00        2,400.00     5,700.00 V
Departmental predetermined OH Rate                8.60              15.40 See U
Overhead applied in Job P     28,380.00      36,960.00 65,340.00 W=V*U
Answer 10- Job Q Molding Fabrication Total
Machine hours used        2,200.00        2,500.00     4,700.00 X
Departmental predetermined OH Rate                8.60              15.40 See U
Overhead applied in Job Q     18,920.00      38,500.00 57,420.00 Y=X*U